OpenStudy (anonymous):

a board game has a spinner with 10 equal sections numbered 1 thr. 10. In order to win the game Mae needs to spin at least a 6 on each of her last two turns. what is the probability of this happening.

OpenStudy (anonymous):

\[\frac{1}{2}\times \frac{1}{2}=\frac{1}{4}\]

OpenStudy (anonymous):

actually it says "at least a 6" so i am assuming that means 6, 7, 8, 9, or 10

OpenStudy (anonymous):

thanks satellite for clarifying. . that would conclude that my interpretation is wrong, lol